# Adapting AI to the Moment: Understanding the Dynamics of Parent–AI Collaboration Modes in Real-Time Conversations with Children

### 29 Mar 2026 v1

@@ 1. Introduction
Parent-AI collaboration in real-time conversational settings with children presents unique challenges. Unlike task-oriented interactions with predefined goals and structures, these settings involve sensitive and open-ended exchanges in which conversational contexts can shift rapidly within a single interaction (Driscoll et al., 2026; Morelen and Suveg, 2012; Lin et al., 2023; Overbeek et al., 2022). As a result, parents’ needs for AI support vary dynamically—for example, they may seek idea generation when fatigued (Dietz Smith et al., 2024), but expect emotional mediation in more sensitive moments (Theofanopoulou and Slovak, 2022). This variability requires parents to continuously decide when and how AI should participate, making it difficult to predefine stable collaboration strategies.
However, existing AI systems for conversational support typically rely on a limited set of static collaboration modes, such as generating prompts or acting as a conversational agent (Ho et al., 2025; Zhang et al., 2022; Liu et al., 2024; Srinivasan and Setlur, 2021). While these approaches demonstrate the potential of AI to assist in conversations, they assume that the set of collaboration modes can be predefined for the entire interaction, providing limited support for adapting AI to continuously evolving conversational contexts.
Despite the growing interest in AI-supported parenting tools, little is known about how parents’ preferred collaboration modes with AI emerge and evolve during ongoing conversations. To address this gap, we investigate how parents configure AI involvement in minimally constrained, real-time interactions. We introduce the concept of collaboration modes to describe how parents configure and engage AI scaffolding functions during interactions.
@@
To answer these questions, we conducted a co-design study with eight parents to elicit parental needs for AI support in parent-child conversations. The study revealed six frequently mentioned functional needs (F1-F6) and two key requirements for collaborating with AI: flexible decision authority and minimal mode-switching costs.
Based on these findings, we developed COMPASS, a collaborative mediation and parental scaffolding system, as a research probe to explore parent-AI collaboration in conversation. COMPASS allows parents to flexibly combine parental support functions during conversations, enabling us to examine how their strategies adapt to shifting collaboration needs.
@@ 2.1. Parent–Child Conversations in HCI
Prior HCI research has explored supporting parent–child conversations from multiple perspectives (Zapf et al., 2023), including (1) content provisioning (e.g., prompts or games) (Chan et al., 2017); (2) process guidance that analyzes conversation dynamics and provides feedback (e.g., tone or emotional cues) (Choi et al., 2025; Dietz Smith et al., 2024); and (3) third-party mediators (e.g., social robots) that restructure interaction (Ho et al., 2024, 2025; Seo et al., 2025). While these approaches provide valuable support, they largely assume stable interaction patterns, limiting adaptation to dynamically evolving conversational needs.
@@ 2.2. AI Systems for Parent-Child Interaction
While these systems explore different role allocations, they primarily focus on static collaboration structures (Li et al., 2025; Wang et al., 2026), with less attention to how parents can adapt AI involvement during ongoing interactions. In practice, parent-child conversations are highly dynamic, with emotions, goals, and engagement shifting in real time (Morelen and Suveg, 2012; Bradley et al., 2015). Supporting parents in responding to these changes requires more in situ control over AI behavior.
@@ 2.3. Collaboration Dynamics in Human-AI Interaction
Recent work instead frames collaboration as “agency negotiation”, where initiative and control shift with task stages and contextual factors (Masson et al., 2024). Adjustable autonomy has been shown to improve perceived usefulness, competence, and satisfaction (Houde et al., 2025; Zhang et al., 2023), and to enhance team performance when AI adapts during collaboration (Salikutluk et al., 2024). While prior work has begun to highlight interaction-level dynamics, most systems examine these dynamics at coarse-grained task units (e.g., storytelling or search sessions) (Zhang et al., 2022; Sharma et al., 2024), with limited work systematically capturing fine-grained interaction changes or validating them through system design.
Recent systems explore finer-grained collaboration mechanisms, including proactive intervention timing (Oh et al., 2024; Kuang et al., 2024; Zhu et al., 2025), proactive task collaboration (Pu et al., 2025; Mei et al., 2025), shared representations for alignment (Jeong et al., 2025; Zhang et al., 2025b), and role fluidity at the micro-task level (Zhou et al., 2024; Liu et al., 2025; Wu et al., 2025).
@@ 3.2. Findings
3.2.1. Scaffolding Functions Aligned with Parental Goals
From the co-design sessions, we identified six frequently requested AI scaffolding functions (F1-F6), which we group into two categories based on Parental Goal Theory (Dix, 1991).
@@ 3.2.2. Flexible Function Combination with Retained Decision Authority
In the co-design sessions, parents emphasized the need to dynamically adapt scaffolding functions to rapidly changing conversational contexts.
@@
Most parents preferred a hybrid approach, where AI suggests an initial set of combinations while leaving final control to parents. P6 noted, “You can offer suggestions for them to adopt,” and P8 suggested “it could prompt whether I need certain functions and I can just select ‘yes’.
Across these views, parents consistently stressed retaining decision authority in choosing functions (P4: “Different parents may have different needs”), as needs vary across situations and conversations are inherently unpredictable (P8).
@@ 4. COMPASS Design and Implementation
Building on insights from the formative study, we designed and implemented COMPASS as a research probe to support flexible parental scaffolding in parent–child conversations. Rather than functioning as a full intervention system, COMPASS is intended to investigate how parents’ collaboration needs and behaviors with AI evolve during such conversations.
To preserve natural decision-making, we minimized system-induced bias that could influence parents’ collaboration choices.
To support analysis of parent-AI collaboration dynamics, we incorporated an annotation interface to capture in-situ reflections and strategies.
@@ 4.1.1. Parent-Child Conversation Interface
Scaffolding functions aligned
@@ 4.1.2. Parent Annotation Interface
For each conversation slice, COMPASS presents the transcript, interface snapshot, and audio replay to support recall. Parents are then asked to respond to structured questions about contextual factors during that slice, including their own state, the child’s state, and interaction-related factors (see all rows labeled “Annotator = Parent” in Table 2). An example of a single annotation is shown in Figure 4.
@@ 5. User Study
We conducted a lab-based study with 21 parent-child pairs using COMPASS as a research probe to examine (1) parents’ preferred AI scaffolding modes, (2) how conversational contextual factors shape these preferences, and (3) strategies parents use to engage with AI during parent–child conversations. The study was IRB-approved, and informed consent/assent was obtained from all participants.
@@ 6.1.1. Moving Towards a More Partner-like collaboration mode
Overall, parents’ preferences shifted toward a more partner-like collaboration mode, characterized by reduced reliance on structural scaffolding (F1, F4) and increased engagement in co-performance with AI (F5, F6).
@@ 6.1.2. Adding Reflective or Expressive Support on Top of Basic Assistance
Parents explored 41 out of 64 possible collaboration modes, indicating diverse and flexible needs beyond fixed configurations. Pairwise correlations among functions were weak ($\phi\in[-0.2,0.2]$; Figure 8), suggesting that functions were combined largely independently.
@@ 6.2.1. Parents’ States Lead to Diverse Collaboration Modes
Cognitive load and energy determine support level. Under high cognitive load, parents prefer direct scaffolding, increasing use of Utterance Suggestion (F2, $p<0.001$) and Conversation Summarization (F4, $p<0.01$). When energy is low, they avoid complex support such as Multimodal Enrichment (F6, $p<0.05$).
@@ 6.3.3. Relationship-Oriented Strategies
Finally, parents used AI to check whether they and their children were “on the same page.” Real-time summaries from F4 reassured them about alignment: “I want to check (whether) our communication is (happening)… (whether) we are engaging in effective communication” (P20).
@@ 7. Discussion
7.1. Evolving Dynamics of Parent-AI Collaboration with COMPASS
Our findings reveal how parent–AI collaboration unfolds over time. We highlight three representative themes that extend prior work and offer new design insights.
7.1.1. From Backstage Support to Partner-style Co-performance
Parents initially relied on backstage support to maintain conversational flow, then gradually granted COMPASS more visible roles (F5, F6), moving toward partner-style collaboration. Extending prior work on flexible role division (Zhang et al., 2022; Gvirsman and Gordon, 2024; Ho et al., 2025), we identify a risk-perception-driven shift: parents incrementally delegate control to assess AI controllability before incorporating it into the “safe zone” of parent–child interaction. This suggests that systems should support progressive trust-building rather than assuming immediate readiness for partnership. It extends HCI trust research into parenting contexts (Niu et al., 2023; Chen et al., 2024b; Jo et al., 2025).
7.1.2. From Experimentation to Stable Collaboration Modes
In the user study, only a subset of collaboration modes was consistently used, while others appeared as short-lived transitional states. Beyond prior work treating collaboration styles as stable preferences (Ho et al., 2024, 2023), our findings show that stable and transitional modes coexist. Transitional modes reflect parents’ experimentation and familiarization with the AI system, echoing exploratory processes in other HCI domains (e.g., exploratory search) (Klouche et al., 2015; Sarma et al., 2024). This suggests that systems should scaffold low-risk experimentation to help families converge on stable strategies suited to their conversational settings.
7.1.3. Context-Aware Restraint in Sensitive Parenting Moments
Parents adapted AI use to context: in calm, cognitively oriented moments, they preferred structured support, whereas in emotionally intense situations, they avoided directive scaffolds and relied on alignment and mediation. While prior HCI work highlights context-aware scaffolding (Ogata and Yano, 2004; Chen et al., 2023, 2025b), we further identify the need for context-aware restraint. AI should not only know when to step in but also hold back in “no-go” moments (e.g., avoiding rigid responses when children are sad). This aligns with “calm technology” (Weiser and Brown, 1996, 1997) and reframes deliberate AI silence as a form of positive intelligence that protects sensitive parent–child interactions.
@@ 7.1.3. Context-Aware Restraint in Sensitive Parenting Moments
7.2. From Parental Goal Theory to AI-Engaged Parental Goal Theory
@@
Parenting goals are typically classified as parent-oriented, child-oriented, and relationship-oriented goals (Dix et al., 1986; Hastings and Grusec, 1998). Grounded in our findings, we revisit these categories and show how they are reshaped in AI-engaged parenting contexts.
7.2.1. Orchestrating AI Support for Parent-Oriented Goals
Prior work shows that parents aim to maintain competence, bridge knowledge gaps, and regulate their own states when interacting with children (Coleman and Karraker, 1998; Deater-Deckard, 1998; Heerman et al., 2017). Our findings support this view, but further show that these goals are increasingly achieved through AI collaboration rather than solely individual effort. Parents treated AI as a just-in-time collaborator that compensated for gaps in knowledge and experience, supported conversational flow, and reduced uncertainty in stressful situations. When facing high cognitive load or emotional instability, they dynamically adjusted collaboration modes to obtain appropriate support, often drawing on familiar technology-use habits. This suggests a shift from individual regulation to parent-AI collaborative regulation in parent-oriented goals.
7.2.2. Extending AI-in-situ Feedback to Support Child-Oriented Goals
Child-oriented goals are typically framed around long-term cognitive and socio-emotional development (Chang and Lee, 2017; Dix, 2014; Ho et al., 2022; Sun et al., 2024), and existing AI systems largely follow this orientation by supporting children’s knowledge learning, emotion recognition, and self-expression (Sun et al., 2024; Chen et al., 2024a, 2025a).
Our findings extend this view by showing that child-oriented goals also involve attending to children’s moment-to-moment states (e.g., emotional type, emotional intensity, and expressive state), which is underexplored in prior theory and system design. While parents are known to regulate children’s emotions during parent-child interactions (Eisenberg and Spinrad, 2004; Eisenberg and Morris, 2002), we find that AI reshapes this process: parents rely on AI to identify and summarize children’s current states, such as emotional type and emotional intensity, reducing the cognitive burden of continuously monitoring children’s states during parent–child conversations.
By activating state-reflective AI features, parents gained a clearer understanding of children’s moment-to-moment states, which directly informed their selection of AI collaboration modes. This suggests that AI expands child-oriented goals from a sole focus on long-term development to a combination of developmental goals and moment-to-moment regulation. Future work should examine how AI can support and balance these two aspects without displacing parental judgment.
7.2.3. AI-Mediated Parent-Child Conversation to Support Relationship-Oriented Goals
Prior work frames relationship-oriented goals as maintaining and enhancing the quality of parent–child relationships (Laursen and Collins, 2009, 2003). Our findings extend this view by showing that AI actively reshapes relational dynamics, as it supports interaction continuity, encourages children’s disclosure, and buffers emotional tension. This is particularly important under high emotional load, where parents often struggle with real-time relationship regulation, leading to escalation and conflict (Barros et al., 2015; Morris et al., 2007).
With AI introduced as a new participant in parent–child conversations, relational maintenance is no longer solely the parent’s responsibility. Parents positioned AI as both a conversational mediator and a relational scaffold: as a neutral third party that encouraged children’s disclosure and helped de-escalate tension, and as a support for alignment and intimacy within the interaction.
This transforms relationship-oriented goals from a dyadic parent–child structure to a triadic parent–child–AI collaboration, extending the theoretical boundary of relational goals. Moreover, parents’ trust and attitudes toward AI shaped how much relational authority they delegated, requiring a balance between AI participation, parental authority, and children’s autonomy. Together, these findings suggest that AI functions as a collaborative social actor in family interaction, fundamentally reshaping the structure of relationship-oriented goals.
7.3. Towards More Fine-grained and Dynamic Human-AI Collaboration Framework
Prior research on human-AI collaboration largely focuses on task-driven contexts, where success is defined by performance metrics such as accuracy and efficiency (Wang and Yin, 2022; Chattopadhyay et al., 2017; Taesiri et al., 2022). Such work often assumes that collaboration goals are clearly defined and relatively stable, and that the division of labor between humans and AI can be specified in advance of interaction. Recent studies introduce human-centered factors (e.g., trust, transparency) to collaboration frameworks (Duan et al., 2025; Vössing et al., 2022), yet often treat these goals in isolation, overlooking how multiple goals interact in practice. Our findings extend this view, showing that parent–AI collaboration in our conversation setting involves navigating multiple, sometimes conflicting goals.
In parent-child conversations, parents simultaneously balance three goals: parent-oriented, child-oriented, and relationship-oriented. The priority of these goals shifts across situations, and parents’ ongoing trade-offs across these goals lead to more fine-grained and dynamic parent-AI collaboration. This goes beyond prior work, which assumes a stable collaboration mode throughout interaction (Ho et al., 2025; Zhang et al., 2022).
In our study, parents switched collaboration modes 7.90 times per conversation on average. With only six functions, they produced 41 of 64 possible combinations, indicating that collaboration modes are not predefined but constructed in situ. Through frequent switching and recombination of AI functions, parents continuously adjust AI’s role, form, and level of involvement in response to conversational progress, the child’s state, and their own judgments. These results show that collaboration modes are emergent, fine-grained, and dynamically reconstructed across time and context.
Together, our findings reframe collaboration success from static performance optimization to the ability to manage multi-objective trade-offs through dynamic adaptation. Parenting represents a high-stakes, relational, and asymmetric-agency context (Gurdal and Sorbring, 2018; Shin et al., 2021; Theofanopoulou and Slovak, 2022). In such contexts, parents must simultaneously assume responsibility while preserving children’s autonomy, which further intensifies the tensions among collaboration goals. Thus, our work also provides a new lens for understanding human-AI collaboration in other high-stakes, relational domains such as healthcare, education, and counseling (Saxena et al., 2021; Xhakaj et al., 2016; Borghoff et al., 2025).

### 01 Apr 2026 v2

@@ 1. Introduction
Parent-AI collaboration in real-time conversational settings with children presents unique challenges. Unlike task-oriented interactions with predefined goals and structures, these settings involve sensitive and open-ended exchanges in which conversational contexts can shift rapidly within a single interaction (Driscoll et al., 2026; Morelen and Suveg, 2012; Lin et al., 2023; Overbeek et al., 2022). As a result, parents’ needs for AI support vary dynamically, ranging from seeking idea generation when fatigued (Dietz Smith et al., 2024) to expecting emotional mediation in sensitive moments (Theofanopoulou and Slovak, 2022). This variability requires parents to continuously decide when and how AI should participate, making it difficult to predefine stable collaboration strategies.
However, existing AI systems for conversational support typically rely on a limited set of static collaboration modes, such as generating prompts or acting as a conversational agent (Ho et al., 2025; Zhang et al., 2022; Liu et al., 2024; Srinivasan and Setlur, 2021). While these approaches demonstrate the potential of AI to assist in conversations, they assume that the set of collaboration modes can be predefined for the entire interaction, providing limited support for adapting AI to evolving conversational contexts.
Despite the growing interest in AI-supported parenting tools, little is known about how parents’ preferred collaboration modes with AI emerge and evolve during ongoing conversations. To address this gap, we investigate how parents configure AI involvement in ongoing interactions. We introduce the concept of collaboration modes to describe how parents configure and engage AI functions during interactions.
@@
To answer these questions, we conducted a co-design study with eight parents to elicit parental needs for AI support in parent-child conversations. We revealed six functional needs (F1-F6) and two key requirements for collaborating with AI: flexible decision authority and minimal mode-switching costs. We then developed COMPASS, a collaborative mediation and parental scaffolding system, as a research probe to explore parent-AI collaboration in conversation. COMPASS allows parents to flexibly combine parental support functions during conversations, enabling us to examine how their strategies adapt to shifting collaboration needs.
@@ 2.1. Parent–Child Conversations in HCI
Prior HCI research has explored supporting parent–child conversations from multiple perspectives (Zapf et al., 2023), including (1) content provisioning (e.g., prompts or games) (Chan et al., 2017); (2) process guidance that analyzes conversation dynamics and provides feedback (e.g., tone or emotional cues) (Choi et al., 2025; Dietz Smith et al., 2024); and (3) third-party mediators (e.g., social robots) that restructure interaction (Ho et al., 2024, 2025; Seo et al., 2025). While these approaches provide valuable support, they largely assume stable interaction, limiting their ability to adapt to dynamic conversations.
@@ 2.2. AI Systems for Parent-Child Interaction
While these systems explore different role allocations, they primarily focus on static collaboration structures (Li et al., 2025b; Wang et al., 2026), with less attention to how parents can adapt AI involvement during ongoing interactions. In practice, parent-child conversations are highly dynamic, with emotions, goals, and engagement shifting in real time (Morelen and Suveg, 2012; Bradley et al., 2015). Supporting parents in responding to these changes requires more in situ control over AI behavior.
@@ 2.3. Collaboration Dynamics in Human-AI Interaction
Recent work frames collaboration as “agency negotiation”, where initiative and control shift with task stages and contextual factors (Masson et al., 2024). Adjustable autonomy can improve perceived usefulness, competence, and satisfaction (Houde et al., 2025; Zhang et al., 2023), and enhance team performance when AI adapts during collaboration (Salikutluk et al., 2024). While prior work highlights interaction-level dynamics, most systems examine these dynamics at coarse-grained task units (e.g., storytelling or search sessions) (Zhang et al., 2022; Sharma et al., 2024), with limited work systematically capturing interaction dynamics or validating them through system design.
Recent systems explore finer-grained collaboration mechanisms, including proactive intervention timing (Oh et al., 2024; Kuang et al., 2024; Zhu et al., 2025), proactive task collaboration (Pu et al., 2025; Mei et al., 2025b), shared representations for alignment (Jeong et al., 2025; Zhang et al., 2025b), and role fluidity at the micro-task level (Zhou et al., 2024; Liu et al., 2025; Wu et al., 2025).
@@ 3.2. Findings
3.2.1. AI Support Functions Aligned with Parental Goals
From the co-design sessions, we identified six frequently requested AI support functions (F1-F6), which we group into two categories based on Parental Goal Theory (Dix, 1991).
@@ 3.2.2. Flexible Function Combination with Retained Decision Authority
In the co-design sessions, parents emphasized the need to dynamically adapt support functions to rapidly changing conversational contexts.
@@
Most parents preferred a hybrid approach, where AI suggests an initial set of combinations while leaving final control to parents. P6 noted, “You can offer suggestions for them to adopt,” and P8 suggested “it could prompt whether I need certain functions and I can just select ‘yes’.” Across these views, parents consistently stressed retaining decision authority in choosing functions, as needs vary across situations and conversations are inherently unpredictable.
@@ 4. COMPASS Design and Implementation
Building on insights from the formative study, we designed and implemented COMPASS. Rather than functioning as a full intervention system, COMPASS is intended to investigate how parents’ collaboration needs and behaviors with AI evolve during such conversations. To preserve natural decision-making, we minimized system-induced bias that could influence parents’ collaboration choices. To support analysis of parent-AI collaboration dynamics, we incorporated an annotation interface to capture in-situ reflections and strategies.
@@ 4.1.1. Parent-Child Conversation Interface
AI support functions aligned
@@ 4.1.2. Parent Annotation Interface
For each conversation slice, COMPASS presents the transcript, snapshot, and audio replay to support recall. Parents are then asked to respond to structured questions about contextual factors during that slice, including their own state, the child’s state, and interaction-related factors (see all rows labeled “Annotator = Parent” in Table 2). An example of a single annotation is shown in Figure 4.
@@ 5. User Study
We conducted a lab-based study with 21 parent-child pairs using COMPASS as a research probe to examine (1) parents’ preferred AI-collaboration modes, (2) how conversational contextual factors shape these preferences, and (3) strategies parents use to engage with AI during parent–child conversations. The study was IRB-approved, and informed consent/assent was obtained from all participants.
@@ 6.1.1. Moving Towards a More Partner-like collaboration mode
Overall, parents’ preferences shifted toward a more partner-like collaboration mode, characterized by reduced reliance on structural support (F1, F4) and increased engagement in co-performance with AI (F5, F6).
@@ 6.1.2. Adding Reflective or Expressive Support on Top of Basic Assistance
Parents explored 41 out of 64 possible collaboration modes, indicating flexible needs beyond fixed configurations. Pairwise correlations among functions were weak ($\phi\in[-0.2,0.2]$; Figure 8), suggesting that functions were combined largely independently.
@@ 6.2.1. Parents’ States Lead to Diverse Collaboration Modes
Cognitive load and energy determine support level. Under high cognitive load, parents prefer direct support, increasing use of Utterance Suggestion (F2, $p<0.001$) and Conversation Summarization (F4, $p<0.01$). When energy is low, they avoid complex support such as Multimodal Enrichment (F6, $p<0.05$).
@@ 6.3.3. Relationship-Oriented Strategies
Finally, parents used AI to check whether they and their children were “on the same page.” Real-time summaries (F4) reassured them about alignment: “I want to check our communication is (happening)… (whether) we are engaging in effective communication” (P20).
@@ 7. Discussion
In this work, we use COMPASS as a research probe to investigate how parents configure AI involvement in real-time conversations with children. Our findings reveal that parent–AI collaboration unfolds as a dynamic, goal-driven, and exploratory process over time. Here, we further discuss (a) how collaboration modes should be conceptualized as adaptive computational entities, (b) how parental goal theory is reshaped in AI-engaged parent–child conversations, and (c) how these insights translate into design implications for parent–AI collaborative systems. Together, these findings inform the design of future family-based systems that enable computable, adaptive collaboration in open-ended educational interactions.
7.1. Toward Computable Adaptive Collaboration in Parent-AI Systems
Prior family-related AI systems have highlighted the importance of flexible AI participation, such as through adjustable parental involvement level or shared role-taking (Zhang et al., 2022; Vargas-Diaz et al., 2023). However, these systems largely operationalize collaboration as a set of predefined roles and participation options.
Our findings extend this line of work by showing that, in sensitive parent-child conversations, collaboration cannot be characterized as a fixed allocation of roles and responsibilities. Instead, parent-AI collaboration modes are continuously recomputed through ongoing trade-offs among multiple parental goals. This dynamic reconfiguration was substantial in our study: parents switched modes an average of 7.90 times per conversation and explored 41 of the 64 possible function combinations. By repeatedly recombining AI functions, parents continuously reconfigured AI’s role, form, and level of involvement in response to evolving conversational contexts.
Besides, this recomputation unfolds through a progressive and exploratory process. Parents typically began with backstage support to maintain conversational control, and gradually granted AI more visible, partner-like roles. This progression was driven by parents’ need to assess AI controllability and safety before incorporating it into the sensitive parent–child interaction space, reflecting a trust-building process. At the same time, not all collaboration modes were equally sustained: while some modes stabilized into recurring patterns, others appeared as short-lived transitional states. These transitional modes reflect parents’ ongoing experimentation and familiarization with AI capabilities, suggesting that collaboration emerges through iterative exploration rather than immediate convergence to fixed strategies.
These findings suggest that, in parent-AI systems, collaboration modes should be explicitly modeled as computational entities that can be continuously inferred and updated. Specifically, collaboration modes can be computed from observable contextual signals (e.g., conversational stage, child state, parental state) to support real-time balancing among competing parental goals, while also adapting to parents’ evolving trust and familiarity with AI. This extends prior HCI work on context-aware and adaptive human-AI interaction systems (Amershi et al., 2019; McGrath et al., 2025).
Building on this insight, we propose a new design direction: family-related interfaces that enable computable adaptive collaboration. Rather than treating collaboration modes as static interface settings, such systems could continuously sense interactional context and adaptively determine how, when, and to what extent AI participates, enabling fluid transitions across collaboration modes. This perspective may also inform the design of human-AI collaboration systems in other sensitive and open-ended HCI domains such as education (Mei et al., 2025a; Chen et al., 2025c; Li et al., 2025a), healthcare (Jo et al., 2025; Xu, 2022), and counseling (Wang et al., 2025; Shah et al., 2022; Hsu et al., 2025).
@@ 7.1. Toward Computable Adaptive Collaboration in Parent-AI Systems
7.2. Revisiting Parental Goal Theory in AI-Engaged Parent-Child Conversation
@@
Parenting goals are typically classified as parent-oriented, child-oriented, and relationship-oriented goals (Dix et al., 1986; Hastings and Grusec, 1998). Grounded in our findings, we revisit these categories and show how they are reshaped in AI-engaged parent-child conversation contexts.
7.2.1. Reconceptualizing Parent-Oriented Goals to Include AI-Mediated Orchestration
Prior work shows that parents aim to maintain competence, bridge knowledge gaps, and regulate their own cognitive and emotional states when interacting with children (Coleman and Karraker, 1998; Deater-Deckard, 1998; Heerman et al., 2017). While our findings support these parent-oriented goals, we extend this view by introducing AI-mediated orchestration: in AI-engaged parent–child conversations, parent-oriented goals are no longer confined to self-regulation, but involve actively configuring AI participation and continuously aligning parents’ mental models of their children.
First, parent-oriented goals expand from self-regulation to the regulation of AI use. Parents act as adaptive orchestrators who continuously determine how, when, and to what extent AI participates in interaction. Rather than using AI as a fixed tool, they iteratively explore system capabilities and dynamically reconfigure collaboration modes, often grounded in their daily technology-use habits. This reframes parent-oriented goals to include managing the AI-mediated interaction process. Moreover, how parents regulate AI use is shaped by their AI literacy (e.g., familiarity and proficiency), which in turn influences how they pursue different levels of AI involvement.
Second, parent-oriented goals extend to the AI-mediated construction of child understanding. Traditionally, parents rely on partial observations filtered through prior beliefs to interpret children’s emotions and behaviors. AI introduces in-situ interpretations of children’s moment-to-moment states (e.g., emotion type, intensity, and expression patterns), which parents appropriate to reflect on and recalibrate their own judgments. In our study, parents actively integrate these AI-generated interpretations into their mental models of their children, which indirectly supports child-oriented and relationship-oriented goals during interaction.
7.2.2. Child-Oriented Goals as Long-Term Development and In-Situ Regulation
While our system does not include AI functions for direct child-orientated support, it supports parents to more effectively pursue their own child-oriented goals within the conversation. Therefore, it remains necessary to examine how AI support reshapes parents’ conceptualizations of child-oriented goals.
Child-oriented goals have traditionally focused on children’s long-term cognitive and socio-emotional development (Chang and Lee, 2017; Dix, 2014; Ho et al., 2022; Sun et al., 2024), with AI systems primarily supporting learning, emotion recognition, and self-expression (Sun et al., 2024; Chen et al., 2024, 2025a). Our findings extend this view by showing that child-oriented goals also involve attending to children’s moment-to-moment states (e.g., emotional type, intensity, and expression), a dimension that remains underexplored in prior work. While parents are known to regulate children’s emotions in interaction (Eisenberg and Spinrad, 2004; Eisenberg and Morris, 2002), AI reshapes this process through real-time monitoring of children’s states. This reframes child-oriented goals as the integration of long-term developmental objectives and in-situ regulation, highlighting a new design space for AI systems to support and dynamically balance these two aspects.
7.2.3. Reconfiguring Relationship-Oriented Goals as Triadic Relationship Orchestration
Prior work frames relationship-oriented goals as maintaining and enhancing the quality of the parent-child relationship (Laursen and Collins, 2009, 2003). Our findings support this view; however, we further reveal that in AI-engaged parent-child conversations, relational goals are no longer confined to dyadic maintenance. Instead, they are reconfigured as the orchestration of a triadic relationship among parent, child, and AI.
AI did not merely provide instrumental assistance during conversation. Parents positioned it as a relational mediator that could sustain interaction, encourage children’s disclosure, and buffer emotional tension—particularly under high emotional load, when parents struggled to regulate conversations without escalating conflict (Barros et al., 2015; Morris et al., 2007). In such moments, AI helped preserve conversational continuity and created a less confrontational interactional space, allowing parents to maintain connection with the child while navigating difficult emotional exchanges.
These findings reconfigure relationship-oriented goals from improving the parent-child bond alone to managing how AI participates in the relationship: when it should act as a neutral mediator, when it should remain in the background, and the degree of relational influence it should be granted. As a result, relationship-oriented goals expand to encompass the coordination of participation, authority, and intimacy across the parent-child-AI triad.
Moreover, parents’ trust in AI and their attitudes toward its social role shape the degree of relational authority they delegate to it. This introduces an ongoing negotiation among AI participation, parental authority, and children’s autonomy. Maintaining a strong parent-child relationship thus increasingly depends on how parents manage AI’s relational role, rather than simply on AI’s presence.